OpenStudy (anonymous):

1/a=b-(1/c), solve for "a"

OpenStudy (kainui):

You could take the inverse of both sides, boom, problem solved.

hero (hero):

\(\frac{1}{a} = b - \frac{1}{c}\) Multiply b by c/c to get: \(\frac{1}{a} = \frac{bc}{c} -\frac{1}{c}\) Combine fractions on the right side to get \(\frac{1}{a} = \frac{bc - 1}{c}\) Cross multiply to get \(c = a(bc-1)\) Divide both sides by bc - 1 to get: \(\large\frac{c}{bc-1} = a\) The end

hero (hero):

Any questions?

OpenStudy (anonymous):

no, that was very helpful, thanks